OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −3) and perpendicular to the line y = 2x + 1? Write your answer as a fraction.

OpenStudy (hoblos):

the product of slopes of two perpendicular lines is -1 m*2 = -1 m=-1/2

Parth (parthkohli):

The slope is the negative reciprocal. The slope here in y = mx + b is m. m here is 2. What is the negative reciprocal of 2?

OpenStudy (anonymous):

So it would just be -2? Or would it be -1/2?

Parth (parthkohli):

As hoblos said...

OpenStudy (anonymous):

Okay, thanks! :)